# Repeating partial disruptions and two-body relaxation

### v1

@@ Abstract
Two-body relaxation may drive stars onto near-radial orbits around a massive black hole, resulting in a tidal disruption event (TDE). In some circumstances, stars are unlikely to undergo a single terminal disruption, but rather to have a sequence of many grazing encounters with the black hole. It has long been unclear what is the physical outcome of this sequence: each of these encounters can only liberate a small amount of stellar mass, but may significantly alter the orbit of the star. We study the phenomenon of repeating partial tidal disruptions (pTDEs) by building a semi-analytical model that accounts for mass loss and tidal excitations. In the empty loss cone regime, where two-body relaxation is weak, we estimate the number of consecutive partial disruption that a star can undergo, on average, before being significantly affected by two-body encounters. We find that in this empty loss cone regime, a star will be destroyed in a sequence of weak pTDEs, possibly explaining the tension between the low observed TDE rate and its higher theoretical estimates.
$*{}^{*}$-mail: ]luca.broggi@unimib.it
@@ 1 Introduction
A massive black hole (MBH) can reveal itself and increase its mass via a so-called stellar tidal disruption event (TDE, Hills(1975); Rees(1988); Phinney(1989); Evans & Kochanek(1989)). A TDE occurs when a star gets too close to the MBH so that the MBH tidal shear overcomes the star self-gravity, resulting in the disruption and accretion of the star by the MBH (e.g. Rees(1988); Lodato et al.(2009)Lodato, King, & Pringle). Such events produce an extremely bright electromagnetic flare observable up to cosmological distances (Bloom et al.(2011)Bloom, Giannios, Metzger, Cenko, Perley, Butler, Tanvir, Levan, O’Brien, Strubbe, De Colle, Ramirez-Ruiz, Lee, Nayakshin, Quataert, King, Cucchiara, Guillochon, Bower, Fruchter, Morgan, & van der Horst; Cenko et al.(2012)Cenko, Krimm, Horesh, Rau, Frail, Kennea, Levan, Holland, Butler, Quimby, Bloom, Filippenko, Gal-Yam, Greiner, Kulkarni, Ofek, Olivares E., Schady, Silverman, Tanvir, & Xu; Leloudas et al.(2016)Leloudas, Fraser, Stone, van Velzen, Jonker, Arcavi, Fremling, Maund, Smartt, Krìhler, Miller-Jones, Vreeswijk, Gal-Yam, Mazzali, De Cia, Howell, Inserra, Patat, de Ugarte Postigo, Yaron, Ashall, Bar, Campbell, Chen, Childress, Elias-Rosa, Harmanen, Hosseinzadeh, Johansson, Kangas, Kankare, Kim, Kuncarayakti, Lyman, Magee, Maguire, Malesani, Mattila, McCully, Nicholl, Prentice, Romero-Cañizales, Schulze, Smith, Sollerman, Sullivan, Tucker, Valenti, Wheeler, & Young) In particular, they can probe the dynamical properties of the galactic nuclei in which they are generated (see e.g. the recent reviews by Stone et al.(2020)Stone, Vasiliev, Kesden, Rossi, Perets, & Amaro-Seoane; Wevers & Ryu(2023)). The most ubiquitous and likely dominant mechanism able to generate a TDE is stellar two-body relaxation: repeated two-body interactions among stars slowly affect the orbital properties of an unlucky star that as a consequence may get too close to the MBH and find itself destroyed. Theoretical works focusing on this aspect typically assess the occurrence rate of TDEs given the properties of the MBH and host environment (e.g. Syer & Ulmer(1999); Magorrian et al.(1998)Magorrian, Tremaine, Richstone, Bender, Bower, Dressler, Faber, Gebhardt, Green, Grillmair, Kormendy, & Lauer; Wang & Merritt(2004)); the growing body of TDE observations (now of the order of $\sim 100$, Gezari(2021)) has sparked growing interest in assessing TDE rates and has recently allowed comparisons between theoretically predicted TDE rates and observationally measured ones. In particular, recent comparisons between theoretical (Stone & Metzger(2016); Stone et al.(2020)Stone, Vasiliev, Kesden, Rossi, Perets, & Amaro-Seoane) and observational (van Velzen(2018); Sazonov et al.(2021)Sazonov, Gilfanov, Medvedev, Yao, Khorunzhev, Semena, Sunyaev, Burenin, Lyapin, Mescheryakov, Uskov, Zaznobin, Postnov, Dodin, Belinski, Cherepashchuk, Eselevich, Dodonov, Grokhovskaya, Kotov, Bikmaev, Zhuchkov, Gumerov, van Velzen, & Kulkarni; Lin et al.(2022)Lin, Jiang, Kong, Huang, Lin, Zhu, & Wang; Yao et al.(2023)Yao, Ravi, Gezari, van Velzen, Lu, Schulze, Somalwar, Kulkarni, Hammerstein, Nicholl, Graham, Perley, Cenko, Stein, Ricarte, Chadayammuri, Quataert, Bellm, Bloom, Dekany, Drake, Groom, Mahabal, Prince, Riddle, Rusholme, Sharma, Sollerman, & Yan) TDE rate estimates suggest that current detection rates are a factor of a few to an order of magnitude lower than empirically-calibrated predictions from classical loss cone theory (though see also Roth et al.(2021)Roth, van Velzen, Cenko, & Mushotzky; Polkas et al.(2023)Polkas, Bonoli, Bortolas, Izquierdo-Villalba, Sesana, Broggi, Hoyer, & Spinoso; Teboul et al.(2024)Teboul, Stone, & Ostriker). This problem may become worse considering the over-representation of observed TDEs in rare post-starburst host galaxies (French et al.(2016)French, Arcavi, & Zabludoff; French et al.(2020)French, Wevers, Law-Smith, Graur, & Zabludoff), pushing down the intrinsic TDE rate in “normal” galaxies.
@@
Theoretical TDE rate calculations usually start from a kinetic theory treatment of the statistical evolution of stellar systems (Chandrasekhar(1942)). Idealizing the stellar distribution as a statistical continuum, one finds that a conical region in velocity space, the “loss cone,” is where TDEs occur, and the rate of TDEs is set by the rate of stellar diffusion into this loss cone (Frank & Rees(1976); Lightman & Shapiro(1977); Shapiro & Marchant(1978); Cohn & Kulsrud(1978)). This theory relies on the assumption that the disruption of a star happens as soon as the pericentre of its orbit is smaller than
@@
known as the tidal disruption radius [^1] A similar approach is also applied to the formation of extreme mass ratio inspirals (EMRIs; Hils & Bender(1995)), which occur when compact objects are captured by MBHs through gravitational wave emission (Hopman & Alexander(2005); Merritt(2013); Broggi et al.(2022)Broggi, Bortolas, Bonetti, Sesana, & Dotti). The EMRI loss cone is not as simple as the classical TDE loss cone, however, because of the slow evolution of EMRI orbits over many pericenter passages; in this way, detailed treatments of EMRI dynamics prefigure the aim of this paper.. Here $m_{\star}$ and $R_{\star}$ respectively represent the stellar mass and radius, while $M_{\bullet}$ is the MBH mass.
@@
In standard loss cone theory, $r_{\rm TDE}$ is typically assumed to be the threshold separation below which the total disruption of the star occurs, while stars approaching with larger pericentres are assumed to remain unperturbed. However, reality is more complex: the star can either be fully destroyed at a pericentre closer than $r_{\rm TDE}$, or, if passing by at larger separations, only some outer layers of the envelope may be "peeled off" the star and accreted by the MBH, while a stellar remnant survives the interaction, giving rise to a partial disruption (e.g. Mainetti et al.(2017)Mainetti, Lupi, Campana, Colpi, Coughlin, Guillochon, & Ramirez-Ruiz; Ryu et al.(2020c)Ryu, Krolik, Piran, & Noble). Such partial disruptions can repeat if the orbit of the star is not dramatically affected by either the partial disruption process itself or by subsequent two-body scatterings over the course of the next orbit. A number of recent works have addressed the expected emission and accretion features associated with the partial disruption of a single star approaching the MBH (Guillochon & Ramirez-Ruiz(2013); Coughlin & Nixon(2019); Miles et al.(2020)Miles, Coughlin, & Nixon), and a growing body of literature is reporting observational candidates for partial disruptions, which manifest themselves as repeating events over timescales of $\sim 0.3-3$ yrs (Payne et al.(2021)Payne, Shappee, Hinkle, Vallely, Kochanek, Holoien, Auchettl, Stanek, Thompson, Neustadt, Tucker, Armstrong, Brimacombe, Cacella, Cornect, Denneau, Fausnaugh, Flewelling, Grupe, Heinze, Lopez, Monard, Prieto, Schneider, Sheppard, Tonry, & Weiland; Wevers et al.(2023)Wevers, Coughlin, Pasham, Guolo, Sun, Wen, Jonker, Zabludoff, Malyali, Arcodia, Liu, Merloni, Rau, Grotova, Short, & Cao; Liu et al.(2023a)Liu, Mockler, Ramirez-Ruiz, Yarza, Law-Smith, Naoz, Melchor, & Rose; Liu et al.(2023b)Liu, Malyali, Krumpe, Homan, Goodwin, Grotova, Kawka, Rau, Merloni, Anderson, Miller-Jones, Markowitz, Ciroi, Di Mille, Schramm, Tang, Buckley, Gromadzki, Jin, & Buchner; Malyali et al.(2023b)Malyali, Liu, Rau, Grotova, Merloni, Goodwin, Anderson, Miller-Jones, Kawka, Arcodia, Buchner, Nandra, Homan, & Krumpe; Malyali et al.(2023a)Malyali, Liu, Merloni, Rau, Buchner, Ciroi, Di Mille, Grotova, Dwelly, Nandra, Salvato, Homan, & Krumpe; Somalwar et al.(2023)Somalwar, Ravi, Yao, Guolo, Graham, Hammerstein, Lu, Nicholl, Sharma, Stein, van Velzen, Bellm, Coughlin, Groom, Masci, & Riddle). However, the theoretical modeling of the event rates associated to partial TDEs is still in its infancy (Stone et al.(2020)Stone, Vasiliev, Kesden, Rossi, Perets, & Amaro-Seoane; Krolik et al.(2020)Krolik, Piran, & Ryu; Chen & Shen(2021); Zhong et al.(2022)Zhong, Li, Berczik, & Spurzem; Bortolas et al.(2023)Bortolas, Ryu, Broggi, & Sesana).
@@
In a recent work, Bortolas et al.(2023)Bortolas, Ryu, Broggi, & Sesana explore for the first time the rate of both partial and total TDEs by extending standard loss cone theory to include partial TDEs, with critical pericenters determined from fully relativistic hydrodynamics simulations involving realistic main sequence stars (Ryu et al.(2020a)Ryu, Krolik, Piran, & Noble; Ryu et al.(2020b)Ryu, Krolik, Piran, & Noble; Ryu et al.(2020d)Ryu, Krolik, Piran, & Noble). In particular, they show that it is possible to relate the rate of total disruptions with the rate of (repeated) partial disruptions that are expected to occur in nuclear clusters; partial disruptions were found to exceed the rate of total disruptions by a factor of ten or more, especially if the stellar system was dominated by the empty loss cone regime (see definition in Sec. 2.2 below). In Bortolas et al.(2023)Bortolas, Ryu, Broggi, & Sesana, the loss cone radius was assumed to be the one at which partial disruptions start to occur, instead of the commonly assumed full disruption radius $r_{\rm TDE}$. However, this first work did not take into account the fact that when a star undergoes multiple partial disruptions, its specific energy can vary owing to the mass stripping it experiences at pericentre, as well as tidal excitation that can be converted into orbital energy.
@@
The paper is organized as follows: in Section 2 we present the physical phenomena included in our model of partial disruptions, and we summarize the key hypotheses and results of classical loss cone theory as presented in Cohn & Kulsrud(1978); in Section 3 we present the results of some simulations that show the effect of repeated disruptions on orbits unaffected by relaxation; finally, in Section 4 we summarize our results and discuss their implications. In Appendix A we show the results of our simulations when accounting for additional effects.
@@ 2 Partial disruptions and loss cone theory
Here we build a simple model of partial disruptions to identify and characterize the key features of the process: the change in stellar mass and specific binding energy. We start by describing how the mass loss experienced by partial disruptions can affect their subsequent orbits.
@@ 2.1 Partial Disruptions
The process of partial disruption around an MBH strongly affects the specific orbital energy $E$ of the subject star, while its angular momentum change is significantly smaller. The orbit of the surviving stellar core is modified by the receipt of a core kick produced by asymmetric mass loss in the envelope; this asymmetry creates a “rocket effect” altering the orbital elements of the core (Manukian et al.(2013)Manukian, Guillochon, Ramirez-Ruiz, & O’Leary). This is in agreement with what found by Ryu et al.(2020c)Ryu, Krolik, Piran, & Noble through hydrodynamical simulations and is consistent with theoretical expectations (see e.g. Ivanov & Novikov(2001)). At the order of magnitude level, the kick imparted to the surviving core, $\delta v_{\rm ml}$, can be written as a fraction $\epsilon$ of the typical velocity at the pericentre $v_{p}\simeq\sqrt{2\,G\,M_{\bullet}/r_{\mathrm{p}}}$, we therefore have
$\frac{\left\lvert\delta E_{\mathrm{ml}}\right\rvert}{E}\simeq\epsilon\,\frac{4\,a}{r_{\mathrm{p}}}\qquad\frac{\delta J_{\mathrm{ml}}}{J}\simeq\epsilon$ (3)
@@
$\frac{\left\rvert\delta E_{\mathrm{tid}}\right\rvert}{E}\simeq\kappa_{E}\,\left(\frac{m_{\star}}{M_{\bullet}}\right)^{2/3}\;\frac{2\,a}{r_{p}}\qquad\frac{\delta J_{\mathrm{tid}}}{J}\simeq\kappa_{J}\,\left(\frac{m_{\star}}{M_{\bullet}}\right)^{2/3}.$ (4)
Therefore, the relative change in orbital angular momentum coming from tidal oscillations may be neglected for the very eccentric orbits we are considering.
@@
where we report the index $n$ of the polytropic model that better agrees with the equilibrium MESA profile (Paxton et al.(2010)Paxton, Bildsten, Dotter, Herwig, Lesaffre, & Timmes) for the two cases.
@@
In the case of model A, we estimate the mass loss as a function of the pericentre through the fitting formulae obtained by Ryu et al.(2020a)Ryu, Krolik, Piran, & Noble; Ryu et al.(2020b)Ryu, Krolik, Piran, & Noble; Ryu et al.(2020c)Ryu, Krolik, Piran, & Noble for stars on nearly radial orbits ($e\gtrsim 0.99$)
@@
that is a smooth extrapolation of the main solution up to the second derivative in the log-log space. Although the simulations of Guillochon & Ramirez-Ruiz(2013) were Newtonian, this is an acceptable approximation for partial disruptions around lower-mass MBHs such as our case. The value of the pericentre where the mass is completely lost gives $\eta^{B}_{\mathrm{TTDE}}=1.11$.
The change in orbital energy due to the ejection of material can be crudely estimated from the asymmetry in the tidal field as (Stone et al.(2013)Stone, Sari, & Loeb; Metzger et al.(2022)Metzger, Stone, & Gilbaum, eq. 32)
@@
The aforementioned equation assumes a pericentre close to the tidal disruption radius $r_{\mathrm{TDE}}$. A more accurate extension to larger pericentres that agrees with hydrodynamical simulations of a star undergoing a partial disruption around an intermediate MBH is simply obtained by replacing $r_{\mathrm{TDE}}$ with the pericentre of the orbit (Kremer et al.(2022)Kremer, Lombardi, Lu, Piro, & Rasio)
@@
An alternative estimate, obtained not from physical reasoning but rather from fitting rational functions to simulations of a star being destroyed by an MBH has been obtained by Gafton et al.(2015)Gafton, Tejeda, Guillochon, Korobkin, & Rosswog
@@
The two prescriptions show significant difference when $\delta m$ is large, while they are comparable (within a factor of 3) when $\delta m\to 0$ (in Fig. 3 they can be compared directly).
As we already mentioned, tidal forces may also alter the stellar structure, inducing oscillations of the star by converting part of its orbital kinetic energy into the mechanical energy of stellar normal modes. This process, known as tidal excitation (or the “dynamical tide”), increases the orbital binding energy in competition with the effects of mass loss. The energy of the oscillations can be written as (Press & Teukolsky(1977))
$\delta E_{\mathrm{tid}}=\left(\frac{G\,M_{\bullet}^{2}}{m_{\star}\,R_{\star}}\right)\,\left(\frac{R_{\star}}{r_{\mathrm{p}}}\right)^{6}\,T(\beta)$ (12)
where $T(\beta)$ is a function that depends on the details of the stellar structure. Assuming that the oscillations are weak, one can treat them in the linear regime, and the function $T(\beta)$ can be expanded, giving
@@
where $T_{2}$, $T_{3}$ are the lowest order terms (quadrupolar and octupolar, respectively) in the linear expansion of the tidal field (Lee & Ostriker(1986)). Although these functions involve complicated integrals over the internal structure of the star, we employ the analytic fitting functions of Portegies Zwart & Meinen(1993), which are accurate for the polytropic models we use.
When mass loss becomes significant and oscillations are strong, the linear regime becomes inadequate; one must instead solve directly the hydrodynamics of the oscillations. While this is most accurately done with hydrodynamical simulations (Manukian et al.(2013)Manukian, Guillochon, Ramirez-Ruiz, & O’Leary), the tidal coupling driving nonlinear oscillations can be approximated reasonably well with the “extended affine model” of Ivanov & Novikov(2001), a generalization of the classical affine model for tidally disrupting stars (Luminet & Carter(1986)) [^3] The affine model approximates the interior of a star subject to strong tidal perturbations as a set of concentric (but not necessarily co-aligned; Ivanov & Novikov(2001)) ellipsoidal shells with time-evolving axis ratios. In this sense, it can be understood as a nonlinear theory of the quadrupolar tide, neglecting higher-order terms.. In the case of some polytrope models, analytic fits of the extended affine model’s predictions for nonlinear $T(\beta)$ have been presented in Generozov et al.(2018)Generozov, Stone, Metzger, & Ostriker. Since stars described by model A are more centrally concentrated with respect to stars of model B – see Fig. 1 – tidal excitation is generally stronger, compared to asymmetric mass loss effects, in Model B ($n=1.5$).
The energy changes presented here allow us to infer whether a star can undergo multiple partial disruptions, and whether it becomes more or less bound at each passage.

### v2

@@ Abstract
Two-body relaxation may drive stars onto near-radial orbits around a massive black hole, resulting in a tidal disruption event (TDE). In some circumstances, stars are unlikely to undergo a single terminal disruption, but rather to have a sequence of many grazing encounters with the black hole. It has long been unclear what is the physical outcome of this sequence: each of these encounters can only liberate a small amount of stellar mass, but may significantly alter the orbit of the star. We study the phenomenon of repeating partial tidal disruptions (pTDEs) by building a semi-analytical model that accounts for mass loss and tidal excitation. In the empty loss cone regime, where two-body relaxation is weak, we estimate the number of consecutive partial disruptions that a star can undergo, on average, before being significantly affected by two-body encounters. We find that in this empty loss cone regime, a star will be destroyed in a sequence of weak pTDEs, possibly explaining the tension between the low observed TDE rate and its higher theoretical estimates.
E-mail: ]luca.broggi@unimib.it
@@ 1 Introduction
A massive black hole (MBH) can reveal itself and increase its mass via a so-called stellar tidal disruption event (TDE, Hills 1975; Rees 1988; Phinney 1989; Evans & Kochanek 1989). A TDE occurs when a star gets too close to the MBH so that the MBH tidal shear overcomes the star self-gravity, resulting in the disruption and accretion of the star by the MBH (e.g. Rees, 1988; Lodato et al., 2009). Such events produce extremely bright electromagnetic flare observable up to cosmological distances (Bloom et al., 2011; Cenko et al., 2012; Leloudas et al., 2016). In particular, they can probe the dynamical properties of the galactic nuclei in which they are generated (see e.g. the recent reviews by Stone et al., 2020; Wevers & Ryu, 2023). The most ubiquitous and likely dominant mechanism able to generate a TDE is stellar two-body relaxation: repeated two-body interactions among stars slowly affect the orbital properties of an unlucky star that as a consequence may get too close to the MBH and find itself destroyed. Theoretical works focusing on this aspect typically assess the occurrence rate of TDEs given the properties of the MBH and host environment (e.g. Syer & Ulmer, 1999; Magorrian et al., 1998; Wang & Merritt, 2004); the growing body of TDE observations (now of the order of $\sim 100$, Gezari, 2021) has sparked growing interest in assessing TDE rates and has recently allowed comparisons between theoretically predicted TDE rates and observationally measured ones. In particular, recent comparisons between theoretical (Stone & Metzger, 2016; Stone et al., 2020) and observational (van Velzen, 2018; Sazonov et al., 2021; Lin et al., 2022; Yao et al., 2023) TDE rate estimates suggest that current detection rates are a factor of a few to an order of magnitude lower than empirically-calibrated predictions from classical loss cone theory (though see also Roth et al. 2021; Polkas et al. 2023; Teboul et al. 2024). This problem may become worse considering the over-representation of observed TDEs in rare post-starburst host galaxies (French et al., 2016, 2020), pushing down the intrinsic TDE rate in “normal” galaxies.
@@
Theoretical TDE rate calculations usually start from a kinetic theory treatment of the statistical evolution of stellar systems (Chandrasekhar, 1942). Idealising the stellar distribution as a statistical continuum, one finds that a conical region in velocity space, the “loss cone,” is where TDEs occur, and the rate of TDEs is set by the rate of stellar diffusion into this loss cone (Frank & Rees, 1976; Lightman & Shapiro, 1977; Shapiro & Marchant, 1978; Cohn & Kulsrud, 1978). This theory relies on the assumption that the disruption of a star happens as soon as the pericentre of its orbit is smaller than
@@
known as the tidal disruption radius [^1] A similar approach is also applied to the formation of extreme mass ratio inspirals (EMRIs; Hils & Bender, 1995), which occur when compact objects are captured by MBHs through gravitational wave emission (Hopman & Alexander, 2005; Merritt, 2013; Broggi et al., 2022). The EMRI loss cone is not as simple as the classical TDE loss cone, however, because of the slow evolution of EMRI orbits over many pericentre passages; in this way, detailed treatments of EMRI dynamics prefigure the aim of this paper.. Here $m_{\star}$ and $R_{\star}$ respectively represent the stellar mass and radius, while $M_{\bullet}$ is the MBH mass.
@@
In standard loss cone theory, $r_{\rm TDE}$ is typically assumed to be the threshold separation below which the total disruption of the star occurs, while stars approaching with larger pericentres are assumed to remain unperturbed. However, reality is more complex: the star can either be fully destroyed at a pericentre closer than $r_{\rm TDE}$, or, if passing by at larger separations, only some outer layers of the envelope may be "peeled off" the star and accreted by the MBH, while a stellar remnant survives the interaction, giving rise to a partial disruption (e.g. Mainetti et al., 2017; Ryu et al., 2020c). Such partial disruptions can repeat if the orbit of the star is not dramatically affected by either the partial disruption process itself or by subsequent two-body scatterings over the course of the next orbit. A number of recent works have addressed the expected emission and accretion features associated with the partial disruption of a single star approaching the MBH (Guillochon & Ramirez-Ruiz, 2013; Coughlin & Nixon, 2019; Miles et al., 2020), and a growing body of literature is reporting observational candidates for partial disruptions, which manifest themselves as repeating events over timescales of $\sim 0.3-3$ yrs (Payne et al., 2021; Wevers et al., 2023; Liu et al., 2023a, b; Malyali et al., 2023b, a; Somalwar et al., 2023). There are a number of hypotheses about the origins of stars on extremely bound orbits, including the tidal break up of stellar binaries (Cufari et al., 2022), the migration of stars through a gaseous disk during the last AGN episode (Metzger et al., 2022), or two-body relaxation in super massive black holes binaries (Melchor et al., 2024). However, the theoretical modeling of the event rates associated to partial TDEs is still in its infancy (Stone et al., 2020; Krolik et al., 2020; Chen & Shen, 2021; Zhong et al., 2022; Bortolas et al., 2023).
@@
In a recent work, Bortolas et al. (2023) explore for the first time the rate of both partial and total TDEs by extending standard loss cone theory to include partial TDEs, with critical pericentres determined from fully relativistic hydrodynamics simulations involving realistic main sequence stars (Ryu et al., 2020a, b, d). In particular, they show that it is possible to relate the rate of total disruptions with the rate of (repeated) partial disruptions that are expected to occur in nuclear clusters; partial disruptions were found to exceed the rate of total disruptions by a factor of ten or more, especially if the stellar system was dominated by the empty loss cone regime (see definition in Sec. 2.2 below). In Bortolas et al. (2023), the loss cone radius was assumed to be the one at which partial disruptions start to occur, instead of the commonly assumed full disruption radius $r_{\rm TDE}$. However, this first work did not take into account the fact that when a star undergoes multiple partial disruptions, its specific energy can vary owing to the mass stripping it experiences at pericentre, as well as tidal excitation that can be converted into orbital energy.
@@
The paper is organised as follows: in Section 2 we present the physical phenomena included in our model of partial disruptions, and we summarise the key hypotheses and results of classical loss cone theory as presented in Cohn & Kulsrud (1978); in Section 3 we present the results of some simulations that show the effect of repeated disruptions on orbits unaffected by relaxation; finally, in Section 4 we summarise our results and discuss their implications. In Appendix A we show the results of our simulations when accounting for additional effects.
@@ 2 Partial disruptions and loss cone theory
Here we build a simple model of partial disruptions to identify and characterise the key features of the process: the change in stellar mass and specific binding energy. We start by describing how the mass loss experienced by partial disruptions can affect their subsequent orbits.
@@ 2.1 Partial Disruptions
The process of partial disruption around an MBH strongly affects the specific orbital energy $E$ of the subject star, while its angular momentum change is significantly smaller. The orbit of the surviving stellar core is modified by the receipt of a core kick produced by asymmetric mass loss in the envelope; this asymmetry creates a “rocket effect” altering the orbital elements of the core (Manukian et al., 2013). This is in agreement with what found by Ryu et al. (2020c) through hydrodynamical simulations and is consistent with theoretical expectations (see e.g. Ivanov & Novikov, 2001). At the order of magnitude level, the magnitude of the kick imparted to the surviving core, $\delta v_{\rm ml}$, can be written as a fraction $\epsilon$ of the typical velocity at the pericentre $v_{p}\simeq\sqrt{2\,G\,M_{\bullet}/r_{\mathrm{p}}}$, we therefore have
$\frac{\left\lvert\delta E_{\mathrm{ml}}\right\rvert}{E}\simeq\epsilon\,\frac{4\,a}{r_{\mathrm{p}}}\qquad\frac{\left\lvert\delta J_{\mathrm{ml}}\right\rvert}{J}\simeq\epsilon$ (3)
@@
$\frac{\left\rvert\delta E_{\mathrm{tid}}\right\rvert}{E}\simeq\kappa_{E}\,\left(\frac{m_{\star}}{M_{\bullet}}\right)^{2/3}\;\frac{2\,a}{r_{p}}\qquad\frac{\left\lvert\delta J_{\mathrm{tid}}\right\rvert}{J}\simeq\kappa_{J}\,\left(\frac{m_{\star}}{M_{\bullet}}\right)^{2/3}.$ (4)
showing that the relative change in orbital angular momentum coming from tidal oscillations may be neglected for the very eccentric orbits we are considering.
@@
where we report the index $n$ of the polytropic model that better agrees with the equilibrium MESA profile (Paxton et al., 2010) for the two cases.
@@
In the case of model A, we estimate the mass loss as a function of the pericentre through the fitting formulae obtained by Ryu et al. (2020a, b, c) for stars on nearly radial orbits ($e\gtrsim 0.99$)
@@
that is a smooth extrapolation of the main solution up to the second derivative in the log-log space. Although the simulations of Guillochon & Ramirez-Ruiz (2013) were Newtonian, this is an acceptable approximation for partial disruptions around lower-mass MBHs such as our case. The value of the pericentre such that $\delta m/m=1$ corresponds to $\eta^{B}_{\mathrm{TTDE}}=1.11$.
The change in orbital energy due to the ejection of material can be crudely estimated from the asymmetry in the tidal field as (Stone et al., 2013; Metzger et al., 2022, Eq. 32)
@@
The aforementioned equation assumes a pericentre close to the tidal disruption radius $r_{\mathrm{TDE}}$. A more accurate extension to larger pericentres that agrees with hydrodynamical simulations of a star undergoing a partial disruption around an intermediate MBH is simply obtained by replacing $r_{\mathrm{TDE}}$ with the pericentre of the orbit (Kremer et al., 2022; Kıroğlu et al., 2023)
@@
An alternative estimate, obtained not from physical reasoning but rather from fitting rational functions to simulations of a star being destroyed by an MBH has been obtained by Gafton et al. (2015)
@@
The two prescriptions differ more when $\delta m$ is large, while they are comparable (within a factor of 3) when $\delta m\to 0$; in Fig. 2 they can be compared directly.
As we already mentioned, tidal forces may also alter the stellar structure, inducing oscillations of the star by converting part of its orbital kinetic energy into the mechanical energy of stellar normal modes. This process, known as tidal excitation (or the “dynamical tide”), increases the orbital binding energy in competition with the effects of mass loss. The specific binding energy converted into oscillations can be written as (Press & Teukolsky, 1977)
$\delta E_{\mathrm{tid}}=\frac{G\,M_{\bullet}^{2}}{m_{\star}\,R_{\star}}\,\left(\frac{R_{\star}}{r_{\mathrm{p}}}\right)^{6}\,T(\beta)$ (12)
where $T(\beta)$ is a dimensionless function that depends on the details of the stellar structure. If the star is already oscillating, whether tidal deformation injects or subtracts energy to the stellar orbit depends on the phase of the ongoing oscillation (Lai, 1997), but Eq. 12 still gives the approximate amplitude of the perturbation. Assuming that the oscillations are weak, one can treat them in the linear regime, and the function $T(\beta)$ can be expanded, giving
@@
where $T_{2}$, $T_{3}$ are the lowest order terms (quadrupolar and octupolar, respectively) in the linear expansion of the star’s response to the tidal field (Lee & Ostriker, 1986). Although these functions involve complicated integrals over the internal structure of the star, we employ the analytic fitting functions of Portegies Zwart & Meinen (1993, Eq. 7 and Tab.1, with $\eta\simeq\beta^{-3/2}$ since $m_{\star}\ll M_{\bullet}$), which are accurate for the polytropic models we use.
When mass loss becomes significant and oscillations are strong, the linear regime becomes inadequate; one must instead solve directly the hydrodynamics of the oscillations. While this is most accurately done with hydrodynamical simulations (Manukian et al., 2013), the tidal coupling driving nonlinear oscillations can be approximated reasonably well with the “extended affine model” of Ivanov & Novikov (2001), a generalisation of the classical affine model for tidally disrupting stars (Luminet & Carter, 1986) [^3] The affine model approximates the interior of a star subject to strong tidal perturbations as a set of concentric (but not necessarily co-aligned; Ivanov & Novikov, 2001) ellipsoidal shells with time-evolving axis ratios. In this sense, it can be understood as a nonlinear theory of the quadrupolar tide, neglecting higher-order terms.. In the case of some polytrope models, analytic fits of the extended affine model’s predictions for nonlinear $T(\beta)$ have been presented in Generozov et al. (2018, Eq. B1 and B2, $\eta\simeq\beta^{-3/2}$).
Figure 2: Magnitudes of energy changes because of asymmetric mass loss (the “rocket effect”) and tidal excitation of oscillation modes for the two stellar models we consider, as a function of the normalized pericentre distance. For each model, we compute the mass lost at a given pericentre according to Eq. 5 (model A) or Eq. 7 (model B) and estimate the corresponding energy change according to Eq. 10 (dashed blue) or Eq. 11 (dashed orange). For a direct comparison, we also plot the magnitude of tidal oscillations according to the non linear theory Eq. 12 (solid green) and the linear expansion (solid pink) Eq. 13. Asymptotically, at large $r_{p}$, the energy change for model A stars is dominated by mass loss, while for model B it is dominated by tidal excitation.
In Fig. 2 we show the competing effects of mass loss and tidal excitation as a function of the pericentre distance for both models. Since stars described by model A are more centrally concentrated with respect to stars of model B – see Fig. 1 – tidal excitation is of greater relative importance in model B ($n=1.5$).
The asymptotic trend at large $r_{p}$ is different fro the two models: model A stars always decrease their orbital binding energy, because of the dominant effect of asymmetric mass loss. In contrast, model B stars can increase or decrease their orbital binding energy, because of the dominant effect of tidal excitation. The energy changes presented here allow us to infer whether a star can undergo multiple partial disruptions, and whether it becomes more or less bound at each passage.